OpenStudy (anonymous):

the legs of an isosceles right triangle are 11cm long. find the length of the hypotenuse. round to the nearest tenth if necessary.

OpenStudy (btaylor):

Use the pythagorean theorem: \[a^2 + b^2 = c^2\] \[c^2 = 11^2 + 11^2 = 2(11^2)\] \[c = \sqrt{2(121)} = 11\sqrt{2} \approx 15.6\]

OpenStudy (anonymous):

thanks